OpenStudy (anonymous):

Please Help!! The supplement of an angle is 10 more than twice the complement of the angle. Find the angle. Thanks!

OpenStudy (anonymous):

Write it in mathlish: 2(90-x)+10=180-x And solve: 180-2x+10=180-x x=10

OpenStudy (anonymous):

Thank you ! I didn't know how to set up this problem